# Semiclassical solution of black hole information paradox

### v2

@@ Abstract
We resolve black hole information paradox within semiclassical gravity, in a manner that does not depend on details of unknown quantum gravity. Our crucial insight is that outgoing Hawking particles are physical only far from the black hole horizon, so they are created far from the horizon and entangled with degrees of freedom closer to the horizon. The later degrees of freedom can be understood as quasi-classical coherent states, implying that Hawking radiation is accompanied with additional radiation similar to classical radiation by which the black hole loses hair during the classical gravitational collapse. The two kinds of radiation are entangled, which resolves black hole information paradox.
@@
Introduction.—Black hole information paradox gid ; math1 ; math2 ; hoss ; dundar ; harlow ; polchinski ; chakra ; marolf ; fabbri is one of the greatest unsolved problems in theoretical physics. The problem appears within semiclassical theory of gravity bd ; mukhanov , which is an incomplete theory where only matter is quantized, while gravity is treated classically. It is widely believed that the correct solution of black hole information paradox depends on details of quantum gravity, while our current understanding of quantum gravity is still very incomplete, thus making the black hole information paradox very difficult to solve. In this paper, however, we argue that the paradox can be solved within semiclassical gravity itself, and that details of quantum gravity are not necessary to understand it. We do not claim that quantum gravity is entirely irrelevant, but we argue that its details are not very important.
@@
One of the central questions that we deal with is where are the outgoing Hawking particles created unruh_where1 ; unruh_where2 ; giddings_where ; hod_where ? The usual picture is that they are created near the horizon. By contrast, we argue that they are created far from the horizon, at a distance much larger than the Schwarzschild radius $2M$ (with $M$ being the mass of the black hole, and we work in units $\hbar=c=G_{N}=k_{B}=1$). Our argument for this claim is based on considerations of foundations of quantum mechanics itself laloe . The contextuality theorems of quantum mechanics KS ; bell ; peres ; laloe show that it is not consistent to assume that measurements just reveal the values of observables that the quantum system possessed before measurements. Instead, it is the measurement that creates physical reality as we perceive it. Of course, the measurement does not need to be performed by a human made apparatus. Instead, the measurement of a quantum subsystem is associated with decoherence decoh1 ; decoh2 caused by its environment, which makes the density matrix of a subsystem diagonal in a certain “preferred” basis determined by interaction with the environment. Applying this logic to Hawking radiation, it follows that it is inconsistent to speak of outgoing Hawking particles before they are measured by the environment. The outgoing Hawking particles are defined as positive frequency modes with respect to the Schwarzschild time coordinate bd , which is a physical time for observers static with respect to the black hole. Such static observers are approximately inertial when they are far from the horizon, namely at $r\gg 2M$. But at smaller distances such observers are non-inertial, so the corresponding “particles” can only be physical if the environment that measures the particles is itself static and thus non-inertial. In principle it is possible to have such a measuring apparatus, but then it measures the particles by the Unruh effect unruh , which requires its own source of energy (such as the fuel of the rocket that keeps a detector hovering at a static position with respect to the black hole), so such a physical realization of Hawking particles at smaller distances is compensated by the loss of energy of this source mochizuki , which does not need to decrease the black hole mass. Thus only particles at $r\gg 2M$ are created by measurement without any additional source of energy for the detector, which makes particles at $r\gg 2M$ much more typical than those at smaller $r$. Hence, for practical purposes, the creation of Hawking particles is naturally interpreted as something that typically happens far from the horizon, at $r\gg 2M$.
@@
The rest of the paper is organized as follows. To build intuition for our approach to black hole information paradox, we first consider two simple analogies, the hydrogen atom information “paradox”, and the case of two entangled harmonic oscillators, in settings that resemble our view of Hawking radiation. Then we consider the quantum state of matter in a classical black hole background and rewrite it in terms of physical Hawking particle states at large distances from the horizon, entangled with degrees of freedom at smaller distances. We rewrite the state of the latter degrees of freedom in the basis of quasi-classical coherent states, and argue that radiation carried by Hawking particles is accompanied with an additional quasi-classical radiation, the latter being similar to radiation produced during the classical gravitational collapse by which a classical black hole loses classical hair. We also discuss the role of more exotic quantum gravitational objects, proposed in the literature as possible solutions of black hole information paradox.
Hydrogen atom information paradox.—Consider a slow electron brought close to a proton. Soon the electron will settle down into the ground state $|\psi_{0}\rangle$ of the hydrogen atom, independently of the initial electron state. This already looks like an apparent loss of information about the initial state, but, of course, the information is not lost, because it is encoded in electromagnetic radiation that electron emits during the process. This is analogous to the creation of a stationary black hole during the gravitational collapse, where, according to the no hair theorem hawking-ellis ; frolov-novikov , most information about the initial state is radiated away in the form of gravitational waves, leaving a stationary black hole characterized by only a few parameters: mass, charge, and angular momentum. The hydrogen atom in the ground state is analogous to the stationary black hole, while information about the initial state of the electron radiated away by electromagnetic radiation is analogous to the initial black hole “hair” radiated away by gravitational waves.
@@
Next suppose that the interaction between the two particles is turned off. The two particles will remain entangled. However, the electron (particle-$A$) will be then captured by the proton again, so it will end up in the ground state $|\psi_{0}\rangle$. The ground state is unique, so its entropy is zero. Particle-$B$, on the other hand, remains in the mixed state with non-zero entropy and it cannot longer be correlated with particle-$A$, because the latter has zero entropy. This establishes the hydrogen atom information “paradox”.
@@
Two harmonic oscillators.—Consider two uncoupled quantum harmonic oscillators, each having the same characteristic frequency $\omega$ and an $n$-basis consisting of states of the usual form
@@
Semiclassical black hole.—Consider a black hole of mass $M$, described either by Schwarzschild $S$-coordinates with coordinate singularity at the horizon at $r=2M$ (where $r$ is the usual Schwarzschild radial coordinate), or by Kruskal $K$-coordinates without a coordinate singularity. The initial state of the quantum field $\phi(x)$ in the classical black hole background is the Kruskal vacuum $|O_{K}\rangle$ bd (known also as Hartle-Hawking vacuum), naturally associated with $K$-coordinates. It is related to the $S$-vacuum $|O_{S}\rangle$, naturally associated with $S$-coordinates, through the formula bd
@@
Our crucial new insight is the following. As we already explained in Introduction, the particles created by $b_{2k}^{\dagger}$ are physical at very large distances $r\gg 2M$, where the spacetime is essentially Minkowski spacetime, but they are not physical at intermediate distances $r$. Hence we split zone-2 into two zones, called zone-$A$ and zone-$B$, defined as follows:
@@
What happens in zone-$(1\!\cup\!A)$?—So far we explained that zone-$(1\!\cup\!A)$, bounded by the surface of large radius $R^{\prime}\gg 2M$, is sufficiently large to accommodate all the entropy needed to resolve the black hole information paradox. Nevertheless, since the particle states in that zone are not physical particles, the question is how to understand that zone in terms of physical objects? In other words, what an inertial observer in that zone would observe?
@@
$|M\rangle=\prod_{k}\int[dZ_{1\cup A}]\sum_{n_{k}=0}^{\infty}\sum_{l_{k}=0}^{n_{k}}\Psi_{n_{k}l_{k}}[Z_{1\cup A}]\;|Z_{1\cup A}\rangle|l_{k}\rangle_{B},$ (27)
@@
$|M\rangle=\prod_{k}\int[dZ_{1\cup A}]\sum_{n_{k}=0}^{\infty}\Psi_{n_{k}n_{k}}[Z_{1\cup A}]\;|Z_{1\cup A}\rangle|n_{k}\rangle_{B}.$ (29)
The $|\Psi_{n_{k}n_{k}}[Z_{1\cup A}]|^{2}$ is the probability density that there will be $n_{k}$ particles (in the $k$-mode) in zone-$B$, and that the fields will be in the quasi-classical coherent state $|Z_{1\cup A}\rangle$ in zone-$(1\!\cup\!A)$.
@@
Quantum gravity exotics?—Our approach, which does not depend on details of quantum gravity, predicts that Hawking radiation is entangled with additional quasi-classical radiation, the latter being similar to classical radiation produced by classical gravitational collapse. In this sense, our approach does not predict any exotic quantum gravitational objects, such as firewalls AMPS , fuzzballs mathur , wormholes er=epr , islands island , baby-universes gid , white holes rovelli_tunnel , supertranslation hair softhair and gravitational crystals nik_cryst , that have been proposed in other approaches to the black hole information paradox. But our approach does not exclude such exotics either. Instead, all such exotic states can be implicitly included in the set of all coherent states contributing to (26). In principle, there may be a non-zero probability for formation of various kinds of such exotic states, depending on details of the unknown quantum theory of gravity. Perhaps even all exotic objects mentioned above have a non-zero probability. But if more than one kind of states $|Z_{1\cup A}\rangle$ contributes, the relevant question is which of these contributions dominates. Such a question cannot be definitely answered without a full theory of quantum gravity. Nevertheless, the most conservative, and perhaps the most plausible, is the scenario in which (29) is dominated by the configurations that maximally resemble the known field configurations in classical general relativity, namely classical hair radiated away by classical mechanisms, as we explained above.
@@
Discussion and conclusion—The crucial ingredient of our approach is the idea that Hawking particles are physical only at large distances $r\gg 2M$ from the black hole, where inertial observers are approximately static with respect to the black hole. At smaller distances it is more physical to describe physics in terms of different objects, that can be entangled with Hawking particles at large distances. Those different objects are most naturally described in terms of quasi-classical fields, leading to the picture in which black hole produces two kinds of radiation, Hawking particles and quasi-classical radiation similar to the radiation produced during the classical black hole collapse. The entanglement between the two kinds of radiation resolves black hole information paradox.
@@
Such resolution of the paradox does not depend on any details of quantum gravity and does not involve any exotic phenomena that would contradict general expectations from semiclassical gravity. In particular, a freely falling observer near the horizon will not observe anything out of the ordinary. Indeed, since Hawking particles are not physical near the horizon, such a freely falling observer near the horizon will not observe any Hawking particles.

### v4

@@ Abstract
We resolve black hole information paradox within semiclassical gravity, in a manner that does not depend on details of unknown quantum gravity. Our crucial insight is that outgoing Hawking particles are physical only far from the black hole horizon, so they are created far from the horizon and entangled with degrees of freedom closer to the horizon. The latter degrees of freedom can be understood as quasi-classical coherent states, implying that Hawking radiation is accompanied with additional radiation similar to classical radiation by which the black hole loses hair during the classical gravitational collapse. The two kinds of radiation are entangled, which resolves black hole information paradox.
@@ I.1 Motivation
Black hole information paradox gid ; math1 ; math2 ; hoss ; dundar ; harlow ; polchinski ; chakra ; marolf ; fabbri is one of the greatest unsolved problems in theoretical physics. The problem appears within semiclassical theory of gravity bd ; mukhanov , which is an incomplete theory where only matter is quantized, while gravity is treated classically. It is widely believed that the correct solution of black hole information paradox depends on details of quantum gravity, while our current understanding of quantum gravity is still very incomplete, thus making the black hole information paradox very difficult to solve. In this paper, however, we argue that the paradox can be solved within semiclassical gravity itself, and that details of quantum gravity are not necessary to understand it. We do not claim that quantum gravity is entirely irrelevant, but we argue that its details are not very important.
I.2 Where are the Hawking particles created?
One of the central questions that we deal with is where are the outgoing Hawking particles created unruh_where1 ; unruh_where2 ; giddings_where ; hod_where ? The usual picture is that they are created near the horizon. By contrast, we argue that they are created far from the horizon, at a distance much larger than the Schwarzschild radius $2M$ (with $M$ being the mass of the black hole, and we work in units $\hbar=c=G_{N}=k_{B}=1$).
Our argument for this claim is based on considerations of foundations of quantum mechanics itself laloe . The contextuality theorems of quantum mechanics KS ; bell ; peres ; laloe show that it is not consistent to assume that measurements just reveal the values of observables that the quantum system possessed before measurements. Instead, it is the measurement that creates physical reality as we perceive it. Of course, the measurement does not need to be performed by a human made apparatus. Instead, the measurement of a quantum subsystem is associated with decoherence decoh1 ; decoh2 caused by its environment, which makes the density matrix of a subsystem diagonal in a certain “preferred” basis determined by interaction with the environment.
@@ I.2 Where are the Hawking particles created?
Applying this logic to Hawking radiation, it follows that it is inconsistent to speak of outgoing Hawking particles before they are measured by the environment. The outgoing Hawking particles are defined as positive frequency modes with respect to the Schwarzschild time coordinate bd , which is a physical time for observers static with respect to the black hole. Such static observers are approximately inertial when they are far from the horizon, namely at $r\gg 2M$. But at smaller distances such observers are non-inertial, so the corresponding “particles” can only be physical if the environment that measures the particles is itself static and thus non-inertial. In principle it is possible to have such a measuring apparatus, but then it measures the particles by the Unruh effect unruh , which requires its own source of energy (such as the fuel of the rocket that keeps a detector hovering at a static position with respect to the black hole), so such a physical realization of Hawking particles at smaller distances is compensated by the loss of energy of this source mochizuki , which does not need to decrease the black hole mass. Thus only particles at $r\gg 2M$ are created by measurement without any additional source of energy for the detector, which makes particles at $r\gg 2M$ much more typical than those at smaller $r$. Hence, for practical purposes, the creation of Hawking particles is naturally interpreted as something that typically happens far from the horizon, at $r\gg 2M$.
I.3 What are the Hawking particles entangled with?
@@
I.4 Organization of the paper
The rest of the paper is organized as follows. To build intuition for our approach to black hole information paradox, in Sec. II we consider two simple analogies, the hydrogen atom information “paradox”, and the case of two entangled harmonic oscillators, in settings that resemble our view of Hawking radiation. Then in Sec. III we consider the quantum state of matter in a classical black hole background and rewrite it in terms of physical Hawking particle states at large distances from the horizon, entangled with degrees of freedom at smaller distances. In Sec. IV we rewrite the state of the latter degrees of freedom in the basis of quasi-classical coherent states, and argue that radiation carried by Hawking particles is accompanied with an additional quasi-classical radiation, the latter being similar to radiation produced during the classical gravitational collapse by which a classical black hole loses classical hair. In Sec. V we discuss the role of more exotic quantum gravitational objects, proposed in the literature as possible solutions of black hole information paradox.
II Simple analogies
II.1 Hydrogen atom information paradox
Consider a slow electron brought close to a proton. Soon the electron will settle down into the ground state $|\psi_{0}\rangle$ of the hydrogen atom, independently of the initial electron state. This already looks like an apparent loss of information about the initial state, but, of course, the information is not lost, because it is encoded in electromagnetic radiation that electron emits during the process. This is analogous to the creation of a stationary black hole during the gravitational collapse, where, according to the no hair theorem hawking-ellis ; frolov-novikov , most information about the initial state is radiated away in the form of gravitational waves, leaving a stationary black hole characterized by only a few parameters: mass, charge, and angular momentum. The hydrogen atom in the ground state is analogous to the stationary black hole, while information about the initial state of the electron radiated away by electromagnetic radiation is analogous to the initial black hole “hair” radiated away by gravitational waves.
@@ II.1 Hydrogen atom information paradox
Next suppose that the interaction between the two particles is turned off. The two particles remain entangled. However, the electron (particle-$A$) is then captured by the proton again, so it ends up in the ground state $|\psi_{0}\rangle$. The ground state is unique, so its entropy is zero. Particle-$B$, on the other hand, remains in the mixed state with non-zero entropy and it cannot longer be correlated with particle-$A$, because the latter has zero entropy. This establishes the hydrogen atom information “paradox”.
@@
II.2 Two harmonic oscillators
Consider two uncoupled quantum harmonic oscillators, each having the same characteristic frequency $\omega$ and an $n$-basis consisting of states of the usual form
@@ II.2 Two harmonic oscillators
III Semiclassical black hole
Consider a black hole of mass $M$, described either by Schwarzschild $S$-coordinates with coordinate singularity at the horizon at $r=2M$ (where $r$ is the usual Schwarzschild radial coordinate), or by Kruskal $K$-coordinates without a coordinate singularity. The initial state of the quantum field $\phi(x)$ in the classical black hole background is the Kruskal vacuum $|O_{K}\rangle$ bd (known also as Hartle-Hawking vacuum), naturally associated with $K$-coordinates. It is related to the $S$-vacuum $|O_{S}\rangle$, naturally associated with $S$-coordinates, through the formula bd
@@ III Semiclassical black hole
Before proceeding, a few notes are in order considering the choices of vacua above. $K$-coordinates and $S$-coordinates are both analogous to Minkowski coordinates in flat spacetime, but in a different sense. $K$-coordinates are analogous to Minkowski coordinates in the sense that they do not contain a coordinate singularity at the horizon. In this sense $S$-coordinates are analogous to Rindler coordinates in flat spacetime, because both contain a coordinate singularity at the horizon. But $S$-coordinates are analogous to Minkowski coordinates in the sense that the components of the metric tensor in Minkowski coordinates look like that in $S$-coordinates at $r\gg 2M$. Hence the assumption that the initial state is the $K$-vacuum corresponds to the assumption that there was no physical Minkowski particles before the black hole was formed. But after the formation of the black hole, the physical particles at $r\gg 2M$ are to be defined with respect to the $S$-vacuum. Of course, this refers to the Schwarzschild black hole, which does not have charge and angular momentum. For more general black holes, with charge and angular momentum, one would need to define analogous vacua with respect to coordinates that generalize $K$-coordinates (by not having coordinate singularities) and $S$-coordinates (by having components of the metric tensor far from the black hole like that in Minkowski coordinates), but in this paper only the simplest Schwarzschild black hole will be considered explicitly.
@@
Our crucial new insight is the following. As we already explained in Sec. I.2, the particles created by $b_{2k}^{\dagger}$ are physical at very large distances $r\gg 2M$, where the spacetime is essentially Minkowski spacetime, but they are not physical at intermediate distances $r$. Hence we split zone-2 into two zones, called zone-$A$ and zone-$B$, defined as follows:
@@
Finally note that, even though $R^{\prime}\gg 2M$, it would not make much sense to take the limit $R^{\prime}\to\infty$. In this limit zone-$A$ would take all of the black hole exterior, so zone-$B$ would not exist. This would mean that physical particles could not exist at any finite distance from the black hole. And this would refer not only to Hawking particles, but to all particles that we usually describe by quantum field theory in flat spacetime, such as those that we observe by standard detectors in particle physics, e.g., at CERN. Infinite $R^{\prime}$ would contradict the fact that we, of course, do observe such particles, despite the existence of black holes at finite distance from Earth. Hence we must render $R^{\prime}$ finite.
IV What happens in zone-$(1\!\cup\!A)$?
So far we explained that zone-$(1\!\cup\!A)$, bounded by the surface of large radius $R^{\prime}\gg 2M$, is sufficiently large to accommodate all the entropy needed to resolve the black hole information paradox. Nevertheless, since the particle states in that zone are not physical particles, the question is how to understand that zone in terms of physical objects? In other words, what an inertial observer in that zone would observe?
@@ IV What happens in zone-$(1\!\cup\!A)$?
$|M\rangle=\int[dZ_{1\cup A}]\prod_{k}\sum_{n_{k}=0}^{\infty}\sum_{l_{k}=0}^{n_{k}}\Psi_{n_{k}l_{k}}[Z_{1\cup A}]\;|Z_{1\cup A}\rangle|l_{k}\rangle_{B},$ (27)
@@
$|M\rangle=\int[dZ_{1\cup A}]\prod_{k}\sum_{n_{k}=0}^{\infty}\Psi_{n_{k}n_{k}}[Z_{1\cup A}]\;|Z_{1\cup A}\rangle|n_{k}\rangle_{B}.$ (29)
The $\prod_{k}|\Psi_{n_{k}n_{k}}[Z_{1\cup A}]|^{2}$ is the probability density that there will be $n_{k}$ particles (in the $k$-modes) in zone-$B$, and that the fields will be in the quasi-classical coherent state $|Z_{1\cup A}\rangle$ in zone-$(1\!\cup\!A)$.
@@ V.1 Quantum gravity exotics?
Our approach, which does not depend on details of quantum gravity, predicts that Hawking radiation is entangled with additional quasi-classical radiation, the latter being similar to classical radiation produced by classical gravitational collapse. In this sense, our approach does not predict any exotic quantum gravitational objects, such as firewalls AMPS , fuzzballs mathur , wormholes er=epr , islands island , baby-universes gid , white holes rovelli_tunnel , supertranslation hair softhair and gravitational crystals nik_cryst , that have been proposed in other approaches to the black hole information paradox. But our approach does not exclude such exotics either. Instead, all such exotic states can be implicitly included in the set of all coherent states contributing to (26). In principle, there may be a non-zero probability for formation of various kinds of such exotic states, depending on details of the unknown quantum theory of gravity. Perhaps even all exotic objects mentioned above have a non-zero probability. But if more than one kind of states $|Z_{1\cup A}\rangle$ contributes, the relevant question is which of these contributions dominates. Such a question cannot be definitely answered without a full theory of quantum gravity. Nevertheless, the most conservative, and perhaps the most plausible, is the scenario in which (29) is dominated by the configurations that maximally resemble the known field configurations in classical general relativity, namely classical hair radiated away by classical mechanisms, as we explained in Sec. IV.
V.2 Tunnelling picture
The description of Hawking radiation in this paper is based on Bogoliubov transformation, which describes the radiation as a state in the Hilbert space and makes the quantum entanglement explicit. However, such a description lacks a simple physical intuitive picture of particle creation. Alternatively, to make particle creation more intuitive, Hawking radiation can be described as quantum tunnelling pw ; zhang . Even though quantum tunnelling in this context is not explicitly formulated in the Hilbert space, so the entanglement is not seen explicitly, a possible resolution of the information paradox can be formulated in terms of energy correlations between particles that tunnel at different times zhang .
In our resolution of the paradox there is no correlation between particles created at different times, but there is a correlation between particles and quasi-classical radiation created at about the same time. In principle, this could also be understood in terms of tunnelling, in a process in which both the radiation and the particles tunnel. Since the quasi-classical radiation is created in zone-$A$, in the tunnelling picture it tunnels through the horizon. By contrast, since the particle is created in zone-$B$, the whole zone-$A$ can be thought of as a classically forbidden zone for the particle, so the particle tunnels not only through the horizon but also through zone-$A$. An explicit technical analysis of such a tunnelling picture is beyond the scope of the present paper, but it can be an interesting challenge for the future work.
V.3 Conclusion
@@
The crucial ingredient of our approach is the idea that Hawking particles are physical only at large distances $r\gg 2M$ from the black hole, where inertial observers are approximately static with respect to the black hole. At smaller distances it is more physical to describe physics in terms of different objects, that can be entangled with Hawking particles at large distances. Those different objects are most naturally described in terms of quasi-classical fields, leading to the picture in which black hole produces two kinds of radiation, Hawking particles and quasi-classical radiation similar to the radiation produced during the classical black hole collapse. The entanglement between the two kinds of radiation resolves black hole information paradox.
@@
Such resolution of the paradox does not depend on any details of quantum gravity and does not involve any exotic phenomena that would contradict general expectations from semiclassical gravity. In particular, the resolution does not involve exotic objects near the horizon, such as firewalls and fuzzballs, so a freely falling observer near the horizon will not observe anything out of the ordinary. Moreover, while in the standard approach, where particles are created near the horizon, one can wonder whether the particle creation will be observed by a freely falling observer, in our approach it is clear that a freely falling observer near the horizon will not observe any Hawking particles, because in our approach Hawking particles are unphysical in the whole zone-$(1\!\cup\!A)$, including the region near the horizon.